The perimeter of a rectangle is 34 m. The length is 8 m more than twice the width. Find the dimensions.
Answer by nerdybill(7384)

The perimeter of a rectangle is 34 m. The length is 8 m more than twice the width. Find the dimensions.
.
Assign variables to things you don't know:
Let W = width
then 2W+8 = length (from:"length is 8 m more than twice the width")
.
Then create your equation:
since we know that the perimeter of any rectangle is:
2(Length+width)
We generate the formula from that:
2((2W+8) + W) = 34
2(3W+8) = 34
6W + 16 = 34
6W = 18
W = 3 m
.
And, we said:
2W+8 = length
2(3)+8 = length
14 m = length
.
Ans:
length = 14 m
width = 3 m